OpenStudy (anonymous):

find number of binomial with x^2-18x+_____=(____)^2

OpenStudy (anonymous):

18/2=9 9^2=81 x^2-18x+81=(x-9)^2

OpenStudy (anonymous):

so its -18;x-18

OpenStudy (anonymous):

no i ment 81 and -9

OpenStudy (anonymous):

you need to complete the square ax^2 + bx + c c= (b/2)^2 in your case (-18/2)^2 = (-9)^2 = 81

OpenStudy (anonymous):

OK i understand it a lot better now thanks a lot :)